OpenStudy (anonymous):

Find the slope and the y-intercept of the line. Write fractional answers in lowest terms. 2y+9x+4=7+9x The slope is ???

OpenStudy (anonymous):

you need to simplify this

OpenStudy (saifoo.khan):

add like terms!

OpenStudy (anonymous):

2y+9x-9x+4 = 7+9x-9x 2y+4 = 7

OpenStudy (anonymous):

=72y+4 = 7+4 =7 No...damn it that's not right.

OpenStudy (anonymous):

I am going to flunk this class. I can see it coming.

OpenStudy (saifoo.khan):

lol.

OpenStudy (anonymous):

It's not funny khan.

OpenStudy (saifoo.khan):

2y+9x+4=7+9x 2y + 4 = 7 +9x -9x 2y + 4 = 7 2y = 7 - 4 2y = 3 y =3/2

OpenStudy (saifoo.khan):

lol, sorry!

OpenStudy (anonymous):

np. YOu have helped me so much

OpenStudy (anonymous):

Okay, so gosh darn it...what is the slope???

OpenStudy (saifoo.khan):

did u got that?

OpenStudy (saifoo.khan):

Lol, slope. the slope is ZERO.

OpenStudy (anonymous):

oh, no I didn't

OpenStudy (anonymous):

You ROCK!!!!

OpenStudy (anonymous):

What would be the y-intercept? (0,?)

OpenStudy (saifoo.khan):

thnks! :) 0,3/2